# Robotic Optimization of Powdered Beverages Leveraging Computer Vision and Bayesian Optimization

### v1

@@ III-D Coffee Optimization
Bayesian Optimization (BO), recognized as one of the most efficient sampling algorithms for black-box functions [24], was selected as the optimization method. It is particularly effective when only a few parameters need to be optimized [47]. BO focuses on creating a surrogate model formed from a Gaussian process that best resembles the examined function. Due to the use of Gaussian processes, this surrogate can also provide a measure of uncertainty, allowing BO to choose points to optimize that reduces uncertainty in the model and find a good solution. The more observations provided to the optimizer, the more confident the algorithm becomes regarding its prediction of optimal parameters. This approach is typically used in cases where evaluation is expensive.
In this study, the indeterministic BO is used as it deemed suitable for sequentially exploring the control parameter design space and reliably quantifying the uncertainty in the system. The goal of the optimization process is to find a set of preparation parameters that minimize the percentage of bubble coverage area visible in the coffee foam to effectively optimize for microfoam. An exploration ratio of $0.6$ was chosen, reflecting the need to explore the design space before exploiting and finding the optimal solution. Despite the efficiency of this approach, searching a design space that has three parameters that exhibit variance and stochastic nature of the results requires tens or hundreds of trials to form an accurate model that can also provide a solution with a good performance.

### v2

@@ III-D Coffee Optimization
Let us consider the objective function $f:\mathcal{X}\rightarrow\mathbb{R}$, which measures the quality of the coffee as the percentage of area without bubbles visible in the coffee foam. $\mathcal{X}$ denotes a bounded domain $\mathcal{X}=[0,60]\times[40,100]\times[10,14]$. A point $\textbf{x}\in\mathcal{X}$ is expressed as $\textbf{x}=(t,s,h)$, where $t$ [s] is mixing time, $s$ [%] is mixing speed, $h$ [cm] is water pouring height. The goal is to maximize the function $f$ over the bounded domain, i.e. we want to find $\arg\max_{\textbf{x}\in\mathcal{X}}f(\textbf{x})$ to effectively optimize for microfoam. As $f(\textbf{x})$ is unknown, Bayesian Optimization (BO), recognized as one of the most efficient sampling algorithms for black-box functions ([24]), was selected as the optimization method. It is particularly effective when only a few parameters need to be optimized ([47]). With BO, a Gaussian Process $\mathcal{GP}$ prior is placed on $f(\textbf{x})$:
$f(\textbf{x})\sim\mathcal{GP}(m(\textbf{x}),k(\textbf{x},\textbf{x}^{\prime})),$
where $m(\textbf{x})$ is the mean function, in this case set to zero, and $k(\textbf{x},\textbf{x}^{\prime})$ is a covariance kernel being a Matern kernel with $\nu=2.5$.
At iteration $n$, $n$ distinct points $\{x^{(i)}\}^{n}_{i=1}\subset\mathcal{X}$ have been observed, with corresponding values $\{y^{(i)}\}^{n}_{i=1}$, where $y^{(i)}=f(\textbf{x}^{(i)})+\epsilon^{(i)}$, with $\epsilon^{(i)}$ accounting for observational noise. Given these data, the posterior predictive distribution for any new point is a normal distribution:
$f(\textbf{x})|\{x^{(i)},y^{(i)}\}^{n}_{i=1}\sim\mathcal{N}(\mu_{n}(\textbf{x}),\sigma^{2}_{n}(\textbf{x})),$
where $\mu$ and $\sigma$ respectively correspond to mean and variance, whose parameters are fitted to the data by maximizing the Gaussian Process’ log marginal likelihood after each observation.
To decide the next experimental point $\textbf{x}_{n+1}$, an Upper Confidence Bound acquisition function was selected:
$\alpha_{n}(\textbf{x})=\mu_{n}(\textbf{x})+\kappa\sigma_{n}(\textbf{x}),$
where $\kappa$ is a parameter for controlling the exploration-exploitation tradeoff. Hence, at step $n$, the next point $\textbf{x}_{n+1}$ is chosen as:
$_{n+1}=\arg\max_{\textbf{x}\in\mathcal{X}}\alpha_{n}(\textbf{x})=\arg\max_{\textbf{x}\in\mathcal{X}}[\mu_{n}(\textbf{x})+\kappa\sigma_{n}(\textbf{x})].$ x ) ] .
The value of $\kappa=8$ was chosen, reflecting the need to explore the design space before exploiting and finding the optimal solution.
The more observations provided to the optimizer, the more confident the algorithm becomes regarding its prediction of optimal parameters. Despite the efficiency of this approach, searching a design space that has three parameters that exhibit variance and stochastic nature of the results requires tens or hundreds of trials to form an accurate model that can also provide a solution with a good performance.